# Designing Elegant Bell Inequalities

### v2

@@ III Main result
We outline the method to derive the condition on the correlation matrix, $F_{n}$, such that maximal quantum value is obtained by generalized EBI correlation. SOS decomposition technique is employed for analytic derivation of maximal quantum value. Consider the shifted operator, ¯$\bar{\mathcal{B}}:=Q\mathds{I}-\mathcal{B}$ and the operators $P_{y}^{n}:=D_{y}^{n}\otimes\mathds{1}-\mathds{1}\otimes(B_{y}^{n})^{\dagger}$ defined with a real constant $Q\in\mathbb{R}$. The matrix $\mathds{I}:=\mathds{1}\otimes\mathds{1}$ is the $d^{2}\times d^{2}$ identity matrix, and $D_{y}^{n}:=\sum_{x}f_{x,y}^{n}A_{x}^{n}$. The SOS decomposition ¯$\bar{\mathcal{B}}=\sum_{n,y}(P_{y}^{n})^{\dagger}P_{y}^{n}\geq 0$ directly implies $\mathcal{B}\right>\leq Q$. Substituting $P_{y}^{n}$ into $\sum_{n,y}(P_{y}^{n})^{\dagger}P_{y}^{n}$ derives
$\mathcal{B}\leq\frac{1}{2}\sum_{n,y}(D_{y}^{n})^{\dagger}D_{y}^{n}\otimes\mathds{1}+\frac{1}{2}\sum_{n,y}\mathds{1}\otimes B_{y}^{n}(B_{y}^{n})^{\dagger}.$ (4)
@@
In the derivation of (4), it is used the Hermiticity of $\mathcal{B}$, i.e. $\mathcal{B}=\sum_{n,y}D_{y}^{n}\otimes B_{y}^{n}=\sum_{n,y}(D_{y}^{n})^{\dagger}\otimes(B_{y}^{n})^{\dagger}$. The decomposition condition, ¯$\bar{\mathcal{B}}=\sum_{n,y}(P_{y}^{n})^{\dagger}P_{y}^{n}$, is obtained by deriving the RHS of (4) as $Q\mathds{I}$. Saturation condition is $\operatorname{tr}[\bar{\mathcal{B}}\rho]=0$ with any valid quantum realization. It implies that $Q$ is attainable upper bound. For all $y,n$, $P_{y}^{n}\rho=0$ is equivalent to the saturation condition [19]. Importantly, this condition is satisfied by the relation $D_{y}^{n}=(B_{y}^{n})^{*}$ on the support of the maximally entangled state as $P_{y}^{n}\left|\phi_{d}^{+}\right>=[D_{y}^{n}\otimes\mathds{1}-\mathds{1}\otimes(B_{y}^{n})^{\dagger}]\left|\phi_{d}^{+}\right>=0$. Lemma $M\otimes\mathds{1}\left|\phi_{d}^{+}\right>=\mathds{1}\otimes M^{\intercal}\left|\phi_{d}^{+}\right>$ is used in the second equation.
@@
Our strategy is to solve the decomposition and saturation condition by varying $f_{x,y}^{n}$ with pre-assumed generalized EBI correlation. To this end, we define the optimal observables satisfying conditions (ii) and (iii),
@@
where $\tau:=\omega^{pq\delta(d,2)/2}$ [39] defined with Kronecker delta $\delta$ and the parameterization $x=dp+q$ is used in Eq.(5). The index $x\in(0,d^{2})$ is now expressed with $p,q\in[0,d)$ satisfying $(p,q)\neq(0,0)$ and $y\in[0,d^{2})$. Alice’s observables in (5) are given as $d^{2}-1$ WH operators. Bob’s observables are generated from $B_{0}$ as Eq.(6). An important part of our method is to construct $B_{0}$ such that it has the fiducial state as its eigenstate. This choice of $B_{0}$ makes it possible to realize WH group covariant SICs on Bob’s eigenspace. Consider $B_{0}=\sum_{\beta}\omega^{\beta}\left|\varphi_{\beta}\right>\left<\varphi_{\beta}\right|$ where the eigenspace is defined as $\Phi(\varphi):=\{\left|\varphi_{\beta}\right>|\beta\in[0,d)\}$, a set of $d$ orthonormal states including the fiducial state $\left|\varphi\right>$. Without the loss of generality, one can let $\left|\varphi\right>:=\left|\varphi_{d-1}\right>$. Then we have,
@@
where the eigenbases of $B_{y}$ are now $\{W_{y}\left|\varphi_{\beta}\right>|\beta\in[0,d)\}$. Consequently, the set of $(d-1)$-th eigenstates of all the Bob’s observables comprises the $d^{2}$ WH group covariant SICs, $\{W_{y}\left|\varphi\right>|y\in[0,d^{2})\}$.
@@
$\tilde{Q}:=\frac{1}{2}\left[\sum_{n}\lVert F_{n}\rVert^{2}+d^{2}(d-1)\right]$ (8)
when $(\mathbf{r}_{a}^{n})^{*}\cdot\mathbf{r}_{x}^{n}=0$ for all cases of $a\neq x$.
Second, we solve the saturation condition with optimal realization, $\left|\phi_{d}^{+}\right>$ and observables (5) and (7). Saturation condition is given as $(B_{y}^{n})^{*}=D_{y}^{n}=\sum_{x}f_{x,y}^{n}W_{x}^{n}=\mathbf{c}_{y}^{n}\cdot\mathbf{A}_{n}$ $\forall y,n$ where $\mathbf{A}_{n}:=\tau^{n}(W_{1}^{n},W_{2}^{n},\ldots,W_{d^{2}-1}^{n})^{\intercal}$. The solution $\mathbf{c}_{y}^{n}$ completely derives the coefficient matrices we after. In sum, we have to solve $(B_{y}^{n})^{*}=\mathbf{c}_{y}^{n}\cdot\mathbf{A}_{n}$ $\forall y,n$ about $\{f_{x,y}^{n}\}$ under constraint $(\mathbf{r}_{a}^{n})^{*}\cdot\mathbf{r}_{x}^{n}=0$ $\forall a\neq x$. The solution is derived in the below theorem.
Theorem. Consider prime $d$, for which WH group covariant SICs is defined. Then, there always exists the expression $S$ defined with,
@@
for $p,q,r,s\in[0,d)$, $dp+q\neq 0$ that satisfies $Q=\tilde{Q}$ with generalized EBI correlation.
@@
Proof.—Consider optimal $B_{0}=\sum_{\beta=0}^{d-1}\omega^{\beta}\left|\varphi_{\beta}\right>\left<\varphi_{\beta}\right|$ defined with eigenspace $\Phi(\varphi)$ containing $\left|\varphi\right>_{d-1}:=\left|\varphi\right>$. This choice of $B_{0}$ realizes SICs in the eigenspace as explained. Saturation condition is re-written with (6) as
$(B_{y}^{n})^{*}$ $=(W_{y}B_{0}^{n}W_{y}^{\dagger})^{*}$
$=\mathbf{c}_{0}^{n}\cdot(W_{y}^{*}\mathbf{A}_{n}W_{y}^{\intercal})$
$=\sum_{p,q}\omega^{-n(ps+qr)}f_{dp+q,0}^{n}(\mathbf{A}_{n})_{dp+q}$
$=\sum_{p,q}f_{dp+q,dr+s}^{n}(\mathbf{A}_{n})_{dp+q}$ (10)
@@
where the notation $(\cdot)_{x}$ stands for the $x$-th element. The second equation is derived from saturation condition, $B_{0}^{n}=(\mathbf{c}_{0}^{n}\cdot\mathbf{A}_{n})^{*}$. Substituting $(\mathbf{A}_{n})_{dp+q}=(\tau X^{p}Z^{q})^{n}$ and $W_{y}=W_{dr+s}=X^{r}Z^{s}$ into RHS of the second equation and suitably applying the canonical commutation relation, $Z^{q}X^{p}=\omega^{pq}X^{p}Z^{q}$, derives the third equation. We used the parameterizations $x=dp+q$ and $y=dr+s$ for the substitution. The equation (10) is obtained from $(B_{dr+s}^{n})^{*}=\mathbf{c}_{dr+s}^{n}\cdot\mathbf{A}_{n}$ and it results in Eq.(9).
@@
Coefficients in Eq.(9) also satisfies the decomposition condition. Substituting (9) into $(\mathbf{r}_{a}^{n})^{*}\cdot\mathbf{r}_{x}^{n}=\sum_{y}(f_{a,y}^{n})^{*}f_{x,y}^{n}$, one can find that it reduces to 0 for all $n$ if $a\neq x$. Finally, showing the saturation condition is always solvable completes the theorem. The elements of $\mathbf{A}_{n}$ is the unitary operator bases for traceless operator [34]. It means that the traceless operator $(B_{0}^{n})^{*}$ can always be expressed as $\mathbf{c}_{0}^{n}\cdot\mathbf{A}_{n}$ with suitable $\mathbf{c}_{0}^{n}$. $\blacksquare$
The above theorem states that the correlation function maximized by generalized EBI correlation always can be found from Eq.(9) with WH group covariant SICs. One has freedom in the choice of SICs, thus $B_{0}$, which determines the coefficients $\{f_{x,0}^{n}\}$ with $(B_{0}^{n})^{*}=\sum_{x}f_{x,0}^{n}(\mathbf{A}_{n})_{x}$. One can construct different correlation functions by varying $\{f_{x,0}^{n}\}$ of Eq.(9) with different choices of SICs. We also note that establishing $\{A_{x}\}$ as unitary operator bases [34] for traceless operator is advantageous for restricting the number of measurements, as the bases are minimum number of unitary operators that express, in linear combination form, an arbitrary traceless unitary operator, in our case $B_{0}^{n}$.
IV Derivation of Bell inequality
IV.1 Gisin’s elegant Bell inequality
@@
IV.2 Bell inequality for two-qutrit system
We derive a Bell inequality maximally violated by generalized EBI correlation in $d=3$. For $d=3$, there are infinitely many fiducial states generating WH covariant SICs, for example $\left|\psi\right>=1/\sqrt{2}(0,1,-\omega^{\theta})$ for $20\leq\theta\leq 1/$ 2 [41]. We choose the one with $\theta=0$, $\left|\varphi\right>:=1/\sqrt{2}(0,1,-1)=1/\sqrt{2}\left(\left|1\right>-\left|2\right>\right)$ which is known to generate the Hesse SICs [42]. With the set of orthonormal bases $\left|\varphi_{0}\right>:=\left|0\right>$, $\left|\varphi_{1}\right>:=1/\sqrt{2}(\left|1\right>+\left|2\right>)$, and $\left|\varphi_{2}\right>:=\left|\varphi\right>$, optimal $B_{0}$ is
@@ IV.2 Bell inequality for two-qutrit system
It is worth mentioning here that our choice of $\left|\varphi_{1}\right>$ is also a fiducial state generating WH covariant SICs. By solving the saturation condition of SOS decomposition, $(B_{0})^{*}=\mathbf{c}_{0}^{1}\cdot\mathbf{A}_{1}$, the first column of $F_{1}$ is easily obtained as $\mathbf{c}_{0}^{1}=(f_{1,0}^{1},f_{2,0}^{1},\ldots,f_{8,0}^{1})^{\intercal}=(1/2)(1,1,\lambda,\nu,\mu,\lambda,\mu,\nu)^{\intercal}$ with $\lambda:=-i/\sqrt{3}$, $\mu:=\omega\lambda$ and $\nu:=\omega^{2}\lambda$. Substituting $\mathbf{c}_{0}$ into Eq.(9) and using the real-valuedness condition $F_{2}=F_{1}^{*}$, the coefficients are fully generated as in [40]. The derived coefficients can be further simplified in a functional form, $f_{dp+q,0}^{n}=1/2$ for $p=0$ and $f_{dp+q,0}^{n}=\omega^{-n(pq+3/2)+3/4}/(2\sqrt{3})$ for $p\neq 0$.
Substituting the derived coefficients into (2), a Bell inequality maximally violated by generalized EBI correlation in two-qurtrit system is derived as,
$S_{3}:=\sum_{n=1}^{2}\sum_{\begin{subarray}{c}p,q,r,s=0\\ 3p+q\neq 0\end{subarray}}^{2}\omega^{-n(ps+qr)}f_{3p+q,0}^{n}\left\langle A_{3p+q}^{n}B_{3r+s}^{n}\right\rangle\leq 15.$ (12)
@@
The upper bound $L=15$ can be obtained in the set of local deterministic correlations satisfying $P(\alpha\beta|xy)=1$ for certain combination of $\alpha,\beta$ and otherwise the probability is 0. The optimal strategy is found as, for all $x$, assigning $P(00|xy)$, $P(01|xy)$, $P(02|xy)$ as 1 respectively for $y=0,1,2$, $y=3,6$, and $y=4,5,7,8$. Violation is obtained from the maximal quantum value, $\tilde{Q}=18$, evaluated from (8) with $\lVert F_{n}\rVert=9$ for $n=1,2$. The upper bound is attained by two parties performing measurements, $A_{x}=W_{x}$ and $B_{y}=W_{y}B_{0}W_{y}^{\dagger}$, on the two-qutrit maximally entangled state when $x\in(0,9)$ and $y\in[0,9)$. As expected, $2d^{2}$ eigenstates, $W_{y}\left|\varphi_{1}\right>$ and $W_{y}\left|\varphi_{2}\right>$ for all $y$, comprise dual sets of $d^{2}$ SICs and eigenstates of four optimal observables for $x=1,3,4,5$, which are given as $Z,X,XZ,XZ^{2}$, defines the complete sets of MUBs [40].
@@
We remark that the Bell inequality defined in (12) shows stronger violation than known Bell inequalities of similar classes. The critical visibility $\nu^{c}$, the smallest visibility $\nu$ of the isotropic state $\rho=\nu\left|\phi_{d}^{+}\right>\left<\phi_{d}^{+}\right|+(1-\nu)\mathds{I}/d^{2}$ to show the violation, is 83.33% for $S_{3}$. It is smaller than that of Bell inequalities including ones defined in $d=3$ for MUBs, 96.77% [19], 96.63% [20], for SICs, 96.41% [20], and original EBI, 86.6%. In addition, the number of settings required for Alice is 8 for $S_{3}$ and it is smaller than the case of known Bell inequality for $d=3$ SICs which requires $\binom{d^{2}}{2}=36$ settings for Alice when the number of Bob’s setting is same as 9 [20].
V Conclusion
We presented a method to construct correlation function for which the quantum maximum is always obtained in a simple analytic form from the generalized EBI correlation. Bell inequality for two-qutrit system derived in our method shows larger violation compared to known Bell inequalities of similar classes, including original EBI. Our work lays the foundation for the future studies for self-testing MUBs and SICs in high dimensional system and DI quantum cryptographic protocols based on it.
Acknowledgements.

### v3

@@ 4 Main result
We present a method to construct Bell expressions, always maximized by generalized EBI correlation if only the WH group covariant SICs is defined in a given dimension. Sum-of-squares (SOS) decomposition technique is used for deriving maximal quantum value in an analytic form. Consider the shifted operator, ¯$\bar{\mathcal{B}}:=Q\mathds{I}-\mathcal{B}$ and operators $P_{y}^{n}:=D_{y}^{n}\otimes\mathds{1}-\mathds{1}\otimes(B_{y}^{n})^{\dagger}$ where $Q\in\mathbb{R}$ is a real constant, $D_{y}^{n}:=\sum_{x}f_{x,y}^{n}A_{x}^{n}$, and $\mathds{1}$ is a $d\times d$ identity matrix. The identity matrix acting on the composite system is denoted by $\mathds{I}:=\mathds{1}\otimes\mathds{1}$. The SOS decomposition ¯$\bar{\mathcal{B}}=\sum_{n,y}(P_{y}^{n})^{\dagger}P_{y}^{n}\geq 0$ directly implies $\mathcal{B}\right>\leq Q$. Substituting $P_{y}^{n}$ into the decomposition derives
$\frac{1}{2}\sum_{n,y}(D_{y}^{n})^{\dagger}D_{y}^{n}\otimes\mathds{1}+\frac{1}{2}\sum_{n,y}\mathds{1}\otimes B_{y}^{n}(B_{y}^{n})^{\dagger}=Q\mathds{I}.$ (7)
@@
In the derivation, the Hermiticity of $\mathcal{B}$, i.e. $\mathcal{B}=\sum_{n,y}D_{y}^{n}\otimes B_{y}^{n}=\sum_{n,y}(D_{y}^{n})^{\dagger}\otimes(B_{y}^{n})^{\dagger}$, is used. Hereafter, the Eq.(7) is to be referred to as ‘decomposition condition’. We define ‘saturation condition’ as $\operatorname{tr}[\bar{\mathcal{B}}\rho]=0$ for any valid quantum realization. It implies that $Q$ is attainable from quantum mechanics. It is known that $P_{y}^{n}\rho=0\forall y,n$ is equivalent to the saturation condition [20]. Importantly, this condition is satisfied by $D_{y}^{n}=(B_{y}^{n})^{*}$ on a maximally entangled state as $P_{y}^{n}\left|\phi_{d}^{+}\right>=[D_{y}^{n}\otimes\mathds{1}-\mathds{1}\otimes(B_{y}^{n})^{\dagger}]\left|\phi_{d}^{+}\right>=[D_{y}^{n}-(B_{y}^{n})^{*}]\otimes\mathds{1}\left|\phi_{d}^{+}\right>$. An identity, $M\otimes\mathds{1}\left|\phi_{d}^{+}\right>=\mathds{1}\otimes M^{\intercal}\left|\phi_{d}^{+}\right>\forall M$, is used in the second equation. This condition reveals the relation between optimal measurements on both subsystems.
@@
Our strategy is to solve the decomposition and saturation conditions simultaneously by varying the coefficients of Bell expression $f_{x,y}^{n}$ with a given generalized EBI correlation. To this end, we define the optimal measurements obeying conditions (ii) and (iii),
@@
where $\tau:=\omega^{pq\delta(d,2)/2}$ [47] is a constant defined with Kronecker delta, $\delta$. The parameterization $x(p,q):=dp+q$ is used in Eq.(8). It allows one to express Alice’s measurements with a single measurement index, $x$. Measurements are specified with $x\in(0,d^{2})$ or $p,q\in[0,d)$ satisfying $(p,q)\neq(0,0)$ and $y\in[0,d^{2})$. Alice’s measurements in (8) are given as $d^{2}-1$ WH operators when $d\geq 3$. Bob’s measurements are generated from $B_{0}$ as Eq.(9). An important part of our method is to construct $B_{0}$ such that it has at least one fiducial state as its eigenstate. This choice of $B_{0}$ makes it possible to realize WH group covariant SICs on Bob’s eigenspace. Consider $B_{0}=\sum_{\beta}\omega^{\beta}\left|\varphi_{\beta}\right>\left<\varphi_{\beta}\right|$ whose eigenspace is defined as $\Phi_{d}:=\{\left|\varphi_{\beta}\right>|\beta\in[0,d)\}$, a set of $d$ orthonormal states including the fiducial state $\left|\varphi\right>$. Without loss of generality, one can let $\left|\varphi\right>:=\left|\varphi_{d-1}\right>$. Then we have,
@@
where $\left|\varphi\right>:=\left|\varphi_{d-1}\right>$. The above spectral decomposition shows that the eigenbases of $B_{y}$ are now $\{W_{y}\left|\varphi_{\beta}\right>|\beta\in[0,d)\}$. Consequently, as $W_{y}\left|\varphi_{d-1}\right>=W_{y}\left|\varphi\right>$, collecting the set of $(d-1)$-th eigenstates of all $d^{2}$ Bob’s measurements, WH group covariant SICs $\{W_{y}\left|\varphi\right>|y\in[0,d^{2})\}$ is obtained. Note that conditions (ii) and (iii) are systematically satisfied with the choice (8) and (10).
@@
$\tilde{Q}(d):=\frac{1}{2}\left[\sum_{n}\lVert F_{n}\rVert^{2}+d^{2}(d-1)\right]$ (11)
when $(\mathbf{r}_{a}^{n})^{*}\cdot\mathbf{r}_{x}^{n}=0$ for $a\neq x$.
Second, we solve the saturation condition with generalized EBI correlation, specifically correlation from $\left|\phi_{d}^{+}\right>$ and measurements (8) and (10). As previously mentioned, the condition can be satisfied with the maximally entangled state and $(B_{y}^{n})^{*}=D_{y}^{n}=\sum_{x}f_{x,y}^{n}A_{x}^{n}=\mathbf{c}_{y}^{n}\cdot\mathbf{A}_{n}$ $\forall y,n$ where $\mathbf{A}_{n}:=(A_{1}^{n},A_{2}^{n},\ldots,A_{d^{2}-1}^{n})^{\intercal}$. From the solution $\mathbf{c}_{y}^{n}$, one obtain a correlation satisfying the saturation condition. In sum, we solve $(B_{y}^{n})^{*}=\mathbf{c}_{y}^{n}\cdot\mathbf{A}_{n}$ $\forall y,n$ about $\mathbf{c}_{y}^{n}$, or about $\{f_{x,y}^{n}\}$, under constraint $(\mathbf{r}_{a}^{n})^{*}\cdot\mathbf{r}_{x}^{n}=0$ $\forall a\neq x$. The derivation is given in the following theorem.
Theorem. Consider prime dimension $d$, in which WH group covariant SICs is defined. Then, there always exists the expression $S$ defined with
@@
for $p,q,r,s\in[0,d)$, $dp+q\neq 0$ satisfying $Q=\tilde{Q}$ with generalized EBI correlation.
@@
Proof.—Consider an optimal measurement $B_{0}=\sum_{\beta=0}^{d-1}\omega^{\beta}\left|\varphi_{\beta}\right>\left<\varphi_{\beta}\right|$ whose eigenspace contains a fiducial state $\left|\varphi\right>_{d-1}:=\left|\varphi\right>$. This choice of $B_{0}$ realizes SICs in the eigenspace, as previously explained. Applying saturation condition for $y=0$, $(B_{0}^{n})^{*}=\mathbf{c}_{0}^{n}\cdot\mathbf{A}_{n}$, into the RHS of Eq.(9), it is derived that
$(W_{y}B_{0}^{n}W_{y}^{\dagger})^{*}$ $=\mathbf{c}_{0}^{n}\cdot(W_{y}^{*}\mathbf{A}_{n}W_{y}^{\intercal})$
$=\sum_{p,q}\omega^{-n(ps+qr)}f_{dp+q,0}^{n}A_{dp+q}^{n}.$ (13)
@@
Substituting $A_{dp+q}^{n}=(\tau X^{p}Z^{q})^{n}$ and $W_{y}=W_{dr+s}=X^{r}Z^{s}$ into the RHS of the first equation and suitably applying the canonical commutation relation, $Z^{q}X^{p}=\omega^{pq}X^{p}Z^{q}$, derives Eq.(13). The parameterization $x=dp+q$ and $y=dr+s$ are used in the derivation. While, the saturation condition for arbitrary $y$, $(B_{y}^{n})^{*}=\mathbf{c}_{y}^{n}\cdot\mathbf{A}_{n}$, reads
$(B_{y}^{n})^{*}$ $=\mathbf{c}_{y}^{n}\cdot\mathbf{A}_{n}$
$=\sum_{x}f_{x,y}^{n}A_{x}^{n}$
$=\sum_{p,q}f_{dp+q,dr+s}^{n}A_{dp+q}^{n}$ (14)
@@
where the same parameterization $x=dp+q$ and $y=dr+s$ are used for comparison with (13). Consequently, one obtain the claimed Eq.(12) from equating (13) and (14). The expression given as (12) also satisfies the decomposition condition. Substituting (12) into $(\mathbf{r}_{a}^{n})^{*}\cdot\mathbf{r}_{x}^{n}=\sum_{y}(f_{a,y}^{n})^{*}f_{x,y}^{n}$, one finds that it reduces to 0 for all $n$ if $a\neq x$. Showing saturation condition is always solvable completes the proof. The elements of $\mathbf{A}_{n}$ is the unitary operator bases for traceless operator [31]. It means that the traceless operator $(B_{0}^{n})^{*}$ can always be expressed as a linear combination form $\mathbf{c}_{0}^{n}\cdot\mathbf{A}_{n}$ with suitable $\mathbf{c}_{0}^{n}$. $\blacksquare$
The above theorem shows that the Bell expression maximized by generalized EBI correlation can always be defined with (12). It is noteworthy that one has freedom in choice of SICs which determine $f_{x,0}^{n}$, the variable parameters in the RHS of (12). With a chosen WH covariant SICs and the corresponding fiducial state, one can construct the optimal $B_{0}$ as previously explained. Then the coefficients for $y=0$, $f_{x,0}^{n}$, are derived from the saturation condition, $(B_{0}^{n})^{*}=\sum_{x}f_{x,0}^{n}(\mathbf{A}_{n})_{x}$. Different Bell expressions can be generated by varying $f_{x,0}^{n}$ in the RHS of Eq.(12) with different choices of SICs.
We also note that assuming Alice’s measurements as unitary operator bases [31] for traceless operator is advantageous for restricting the number of measurements in our method. The traceless unitary operator bases $\{U_{x}\}$ are the minimum number of unitary operators that satisfy $V=\sum_{x}v_{x}U_{x}$ for an arbitrary traceless unitary operator $V$. Therefore, with our choice of Alice’s measurements, $B_{0}$ constructed from any fiducial vector can be considered to satisfy the saturation condition, $(B_{0}^{n})^{*}=\sum_{x}f_{x,0}^{n}(\mathbf{A}_{n})_{x}$. In other words, any WH covariant SICs can be considered. As a Bell test requires random inputs to close freedom-of-choice loophole [50], reducing the number of settings can be advantageous for lowering input randomness. Also, fewer measurements might beneficial to restrict the number of experimental parameters to be controlled.
5 Derivation of Bell inequality
5.1 Gisin’s elegant Bell inequality
@@
5.2 Bell inequality for two-qutrit system
We derive a Bell inequality maximally violated by generalized EBI correlation in $d=3$. For $d=3$, there are infinitely many fiducial states generating WH covariant SICs, for example $\left|\psi\right>=1/\sqrt{2}(0,1,-\omega^{\theta})$ for $20\leq\theta\leq 1/$ 2 [49]. We choose one with $\theta=0$, $\left|\varphi\right>=1/\sqrt{2}(0,1,-1)=1/\sqrt{2}\left(\left|1\right>-\left|2\right>\right)$ which is known to generate the Hesse SICs [42]. With the set of orthonormal bases $\left|\varphi_{0}\right>:=\left|0\right>$, $\left|\varphi_{1}\right>:=1/\sqrt{2}(\left|1\right>+\left|2\right>)$, and $\left|\varphi_{2}\right>:=\left|\varphi\right>$, optimal $B_{0}$ is constructed as
@@ 5.2 Bell inequality for two-qutrit system
It is worth mentioning here that our choice of $\left|\varphi_{1}\right>$ is also a fiducial state for another WH covariant SICs. By solving the saturation condition, $(B_{0})^{*}=\mathbf{c}_{0}^{1}\cdot\mathbf{A}_{1}$, the first column of $F_{1}$ is straightforwardly obtained as $\mathbf{c}_{0}^{1}=(f_{1,0}^{1},f_{2,0}^{1},\ldots,f_{8,0}^{1})^{\intercal}=(1/2)(1,1,\lambda,\nu,\mu,\lambda,\mu,\nu)^{\intercal}$ with $\lambda:=-i/\sqrt{3}$, $\mu:=\omega\lambda$ and $\nu:=\omega^{2}\lambda$. Substituting $\mathbf{c}_{0}$ into Eq.(12) and using the real-valuedness condition $F_{2}=F_{1}^{*}$, the coefficients are fully generated as presented in [48]. The derived coefficients for the first column can be further simplified in a functional form,
$f_{dp+q,0}^{n}=\begin{cases}\frac{1}{2},\quad(p=0)\\ \frac{i}{2\sqrt{3}}(-1)^{n}\omega^{-npq},\quad(p\neq 0).\end{cases}$ (16)
Substituting the derived coefficients (16) into (12), a Bell expression is defined as:
$S_{3}:=\sum_{n=1}^{2}\sum_{\begin{subarray}{c}p,q,r,s=0\\ 3p+q\neq 0\end{subarray}}^{2}\omega^{-n(ps+qr)}f_{3p+q,0}^{n}\left\langle A_{3p+q}^{n}B_{3r+s}^{n}\right\rangle.$ (17)
A Bell inequality for generalized EBI correlation in two-qutrit system is derived as $S_{3}\leq 15$. The upper bound $L=15$ can be obtained in the set of local deterministic correlations obeying $P(\alpha\beta|xy)=1$ for a certain combination of $\alpha,\beta$ and otherwise the probability is 0. The optimal strategy is found as, for any $x$ and respectively for $y=0,1,2$, $y=3,6$, and $y=4,5,7,8$, assigning $P(00|xy)$, $P(01|xy)$, $P(02|xy)$ as 1.
@@
Violation is obtained from the maximal quantum value, $\tilde{Q}=18$, evaluated from (11) with $\lVert F_{1}\rVert=\lVert F_{2}\rVert=9$. It is obtained by two parties performing measurements, $A_{x}=W_{x}$ and $B_{y}=W_{y}B_{0}W_{y}^{\dagger}$, on a two-qutrit maximally entangled state $|\phi^{+}_{3}\rangle$ when $x\in(0,9)$ and $y\in[0,9)$. As we considered two fiducial states, $\left|\varphi_{1}\right>$ and $\left|\varphi_{2}\right>$, $2d^{2}$ eigenstates on Bob’s side $\{W_{y}\left|\varphi_{1}\right>|y\in[0,9)\}$ and $\{W_{y}\left|\varphi_{2}\right>|y\in[0,9)\}$ make dual SICs. The eigenstates of Alice’s four optimal observables for $x=1,3,4,5$, namely $Z,X,XZ,XZ^{2}$, define the complete MUBs [48].
@@
We remark that the Bell inequality $S_{3}$ shows stronger violation than known Bell inequalities of similar classes. The critical visibility $\nu^{c}$, the smallest visibility $\nu$ of the isotropic state $\rho=\nu|\phi_{d}^{+}\rangle\langle\phi_{d}^{+}|+(1-\nu)\mathds{I}/d^{2}$ to show the violation, is 83.33% for $S_{3}$. It is smaller than that of Bell inequalities defined for $d=3$ MUBs, 96.77% [20], 96.63% [21], SICs, 96.41% [21], and original EBI, 86.6%. In addition, the number of settings required for Alice is 8 for $S_{3}$ and it is smaller than the case of known Bell inequality for $d=3$ SICs which requires $\binom{d^{2}}{2}=36$ settings for Alice when the number of Bob’s setting is same as 9 [21].
6 Conclusion
We present a method to construct Bell expressions for which the quantum maximum is always obtained in a simple analytic form from the generalized elegant Bell inequality (EBI) correlation. That any WH SICs can be considered for the construction provides versatile nature for our scheme. It is also discussed that the number of measurement settings are restricted while maintaining the versatility.
For example, we derived the original EBI and a novel Bell inequality for generalized EBI correlation by applying our method respectively for two and three dimension. It is notable that the derived Bell inequality shows larger violation compared to known Bell inequalities, either for three dimensional MUBs or SICs. It improves the visibility of isotropic state more than 13%.
Our primary contribution lies in suggesting a method to derive Bell expressions maximized by generalized EBI correlation. The exploration of Bell inequalities in arbitrary higher dimensions represents an interesting direction for future research, requiring broader analysis such as finding suitable choice of SICs in high dimension. We believe our work provides a key framework for such exploration. This work also naturally establishes a foundation for future research on self-testing MUBs and SICs in high-dimensional systems, as well as DI quantum cryptographic protocols which would be built upon them.